OpenStudy (anonymous):

PLEASE HELP!! :( I have this y=mx+B form -> y=-0.031ln(x) + 0.1306 have to plug in 2 for x. so it would be y=-0.031ln(2) + 0.1306 how would I solve this? I forgot the log rules and I need this for a bio lab report. please help me. Thank you so much!

OpenStudy (anonymous):

should I find the Ln of 2 first than multiply the -0.031 and add 0.1306?....

jimthompson5910 (jim_thompson5910):

y=-0.031ln(x) + 0.1306 is not in the form y = mx+b though

OpenStudy (anonymous):

yeah it is.well, technically it won't seem like it but yes. it is. the m=-0.031 x=2 b=0.1306

jimthompson5910 (jim_thompson5910):

alright so if m = -0.031 and b = 0.1306, then y = mx+b y = -0.031x+0.1306

jimthompson5910 (jim_thompson5910):

assuming it is in y=mx+b form

OpenStudy (gorv):

ln 2=0.693

OpenStudy (gorv):

just plug in this value in place of ln2 and do the rest of operation required and find y

OpenStudy (anonymous):

so I find the Ln first than just do the rest?

OpenStudy (anonymous):

Well you can always download a graphing utility (graphcalc, or some other scientific calculator) with an (ln) operator and replace y with f(x), so you would be working with a linear function when, x = 2. f(x) = -0.031ln(x) + 0.1306 f(2) = -0.031ln(2) + 0.1306 And then evaluate when x = 2, at f(2) = ? Type in the calculator ln(2), multiply by -0.031, then add 0.1306

OpenStudy (anonymous):

I was trying to do it with excel but it was weird and didn't give me an answer. so I have to solve it by hand. I just didn't know where the Ln was in regards to the order of operations

OpenStudy (anonymous):

Also, considering this is in slope-intercept form, you can graph it relatively simple as well. Here is the natural log graph, what is y when x = 2? That is the same question as what is f(2) in the function f(x), when f(x) = -0.031ln(x) + 0.1306

OpenStudy (anonymous):

Here is the graph with the cursor just about where x = 2. This is the graphical representation of the solution f(2) to the linear equation shown algebraically as f(x) = -0.031ln(2) + 0.1306

OpenStudy (anonymous):

yeah but I need to find the value for this, another X and divide by 2. that would be slope 1...well for output 1 in this report. thanks for the graph! I actually downloaded that graph software. I plugged in the f(x) equation as you told me, and it's not solving. I'll just do everything in my calc. but what did u get y as ? for the original q after solving everything? I got 0.1090 (after rounding)